You are given the following information: Stockholders' equity as reported on the firm’s balance sheet = $5.75 billion, price/earnings ratio = 20.5, common shares outstanding = 180 million, and market/book ratio = 1.4. The firm's market value of total debt is $7 billion, the firm has cash and equivalents totaling $350 million, and the firm's EBITDA equals $1 billion. What is the price of a share of the company's common stock? Do not round intermediate calculations. Round your answer to the nearest cent.

What is the firm's EV/EBITDA? Do not round intermediate calculations. Round your answer to two decimal places.
Answer #1

- Price/Earning ratio= 20.5

MPS/EPS = 20.5

MPS = 20.5*EPS

- Market Value = Market price(MPS)*No. of share outstanding

Market/Book ratio = 1.4

20.5*EPS*180 million/5750 million = 1.4

3690*EPS = 8050

EPS = 8050/3690

EPS =$ 2.1816

Taking EPS in "MPS=20.5*EPS"

MPS = 20.5*2.1816

MPS = $ 44.72

So, price of a share of the company's common stock is $ 44.72

Calculating Firm's Enterprise Value(EV):

EV = Market capitalization + Preferred Stocks + Outstanding Debt +Minority Interest - Cash & Cash Equivalents

= (44.72*180 million) + 0 + 7000 million + 0 - 350 million

= $ 14699.6 million

EBITDA = $ 1000 million

Firm's EV/EBITDA = $ 14699.6 million/$ 1000 million

= 14.70 times

Note- If you need any clarification regarding this solution, then you can ask in comments

If you like my answer then please Up-vote as it will be motivating.